40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3108     Entry time: Wed Jun 23 17:48:16 2010
Author: steve 
Type: Update 
Category: MOPA 
Subject: laser head temp 

The laser chiller temp is fluctuating and the power output is decreasing. See 120 days plot.

Yesterday I removed ~300cc water from the overflowing chiller tank.

Attachment 1: htemp120d.jpg  86 kB  | Hide | Hide all
htemp120d.jpg
ELOG V3.1.3-